# Quantum resonances and scattering poles of classical rank one locally symmetric spaces

### v1

@@ 1. Introduction
The scattering poles mentioned in the title are the poles of the meromorphic families $S_{\lambda}$ and $S_{\lambda}^{\Gamma}$ of scattering matrices which are constructed from the standard intertwining operators for the spherical principal series representations of $G$. For the definition of the scattering matrices $S_{\lambda}^{\Gamma}$ on $\Gamma\backslash X$ we build on the constructions from [BO00], which for technical reasons exclude the case $G=F_{4}^{-20}$. This is the only reason why we restrict our attention to the classical rank one spaces.
@@
The relation between quantum resonances and scattering poles for the global rank one symmetric spaces was described explicitly in [HHP19]. The quantum resonances are given as the zeros of the function $c(i\zeta)c(-i\zeta)$ with $\mathrm{Im}\,\zeta>0$, where $c$ is a variant of the Harish-Chandra $c$-function. The zeros of $c$ are known (see e.g. [HHP19, Rem. 5.1]). [HHP19, Thm. 7.1]. The main result of that paper says that the scattering poles contained in $\mathcal{H}_{\mathrm{phys}}$ are precisely the quantum resonances, whereas the scattering poles which are not quantum resonances are the poles of the family of standard intertwining operators.
@@
In [Hil24, Problem 6.5] it was suggested that one can use results of Bunke and Olbrich to extend the relation between quantum resonances and scatttering poles to convex-cocompact locally symmetric spaces of rank one. In fact, in [BO12] Bunke and Olbrich studied the meromorphic continuation of the resolvent kernels as well as the scattering poles for the classical locally symmetric spaces with negative curvature and convex-cocompact fundamental group. Their analysis is based on the meromorphic continuation of a family of operators extending $\Gamma$-invariant distributions on $\Omega$ to $\Gamma$-invariant distributions on $\partial X$. Our point of departure is [BO12, Lemma 11], where they show that the difference between the resolvent kernels on $X$ and the lift of the resolvent kernels on $\Gamma\backslash X$ has boundary values on $\partial X\times\partial X$ which depend meromorphically on the spectral parameters. A careful analysis of the proof of this lemma yields a meromorphic function relating the the boundary values of the lifted resolvent kernels of $\Gamma\backslash X$ and the lifted kernels of the scattering matrices. This allows us to generalize the statement about the scattering poles being quantum resonances except for the exceptional spectral parameters for which the Poisson transform is singular so that no boundary value map is defined. More precisely, we can define a physical half-plane $\mathcal{H}^{\Gamma}_{\mathrm{phys}}$ depending on $\Gamma$ such that our main result can be formulated as follows (see Theorem 5.4).
@@
Theorem. Outside the set $-\frac{i}{2}\left\|\alpha_{0}\right\|\mathbb{N}$, the scattering poles which are not in the physical plane $\mathcal{H}^{\Gamma}_{\mathrm{phys}}$ are precisely the quantum resonances.
Here $\alpha_{0}$ is the unique reduced root in a chosen positive system, see Section 2.
@@ Acknowledgements
BD has received funding from the Deutsche Forschungsgemeinschaft (German Research Foundation, DFG) through the Priority Program (SPP) 2026 “Geometry at Infinity”. JH acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via the grant SFB-TRR 358/1 2023 — 491392403.
@@ 3. Resolvent kernels
$\mathcal{H}^{\Gamma}_{\mathrm{phys}}:=\begin{cases}\{\zeta\in\mathbb{C}\mid\mathrm{Im}\,\zeta>0\}&\text{for }0\leq\delta_{\Gamma}\leq\|\rho\|\\ \{\zeta\in\mathbb{C}\mid\mathrm{Im}\,\zeta>\delta_{\Gamma}-\|\rho\|\}&\text{for }\|\rho\|<\delta_{\Gamma}\leq 2\|\rho\|.\end{cases}$
@@
$c_{\Gamma}:=\delta_{\Gamma}^{2}+\max(0,\delta_{\Gamma}-\left\|\rho\right\|)^{2}\in[\delta_{\Gamma}^{2},2\delta_{\Gamma}^{2}].$ (3.3)
@@
$\big{|}((e,\gamma)^{\ast}r_{\zeta})(f)\big{|}$ $\leq C\left\|f\right\|_{\infty}e^{-a\sqrt{-\max(0,\delta_{\Gamma}-\left\|\rho\right\|)^{2}-\mathrm{Re}\,(\zeta^{2})}}\quad\forall\;\zeta\in W$
@@ 4.1. Spherical principal series representations
$\sigma_{\lambda}:P=MAN\to\mathbb{C}\setminus\{0\},\quad man\mapsto a^{\rho+\lambda},$ (4.1)
@@
is an isomorphism of vector spaces. This allows us to consider $H^{\lambda,\infty}$ and $C^{\infty}(G/P;V(\sigma_{\lambda}))$ as $\lambda$-independent vector spaces. It also shows that $V(\sigma_{\lambda})$ is in fact a trivial line bundle but we will soon pass to its $\Gamma$-quotient, which need no longer be trivial.
@@
over the smooth manifold $\Gamma\backslash\Omega$. The vector space $C^{\infty}(\Gamma\backslash\Omega;V_{\Gamma\backslash\Omega}(\sigma_{\lambda}))$ corresponds canonically to the space of $\Gamma$-invariant smooth sections of $V(\sigma_{\lambda})|_{\Omega}$, which in turn corresponds canonically to the closed subspace of left-$\Gamma$-invariant functions in $H^{\lambda,\infty}$. Thus, the space $C^{\infty}(\Gamma\backslash\Omega;V_{\Gamma\backslash\Omega}(\sigma_{\lambda}))$ inherits a Fréchet topology from $H^{\lambda,\infty}$. In particular, we can consider $C^{\infty}(\Gamma\backslash\Omega;V_{\Gamma\backslash\Omega}(\sigma_{\lambda}))$ as a $\lambda$-independent Fréchet space, too.
@@ 4.2. Comparing conventions
Remark 4.2.
Our definitions (4.1), (4.2) differ from the one in [BO00, Section 3], [Olb94, § 2.2] and [HHP19, § 2] by the sign of $\lambda$. We chose this sign convention to be consistent with [Kna86, VII, §1]. This is also consistent with the sign convention in [Vog08, Chap. 3]. In all these references the $G$-action is the left regular action. On the other hand, [Wal92, § 10.1] and [KS71] use the right regular version of the principal series representations. Note that the map $f\mapsto\tilde{f}$ with $\tilde{f}(g):=f(g^{-1})$ intertwines left and right regular actions. Up to this intertwiner the sign conventions in [Wal92, § 10.1] and [KS71] agree with the sign conventions in [Kna86, VII, §1]. Finally, we point out that our topologies on $C^{\infty}(G/P;V(\sigma_{\lambda}))$ and $C^{\infty}(\Gamma\backslash\Omega;V_{\Gamma\backslash\Omega}(\sigma_{\lambda}))$ are equivalent to those of [BO00, p. 89].
Remark 4.3 (Compatibility with [BO12, Section 6] and [HHP19]).
@@
~±${\widetilde{\mathbb{C}}}^{\pm}\owns z\mapsto\sqrt{\pm(\left\|\rho\right\|^{2}-z)}\in\mathbb{C}$ (4.6)
are holomorphic, respectively. As Riemann surfaces, both ${\widetilde{\mathbb{C}}}^{+}$ and ${\widetilde{\mathbb{C}}}^{-}$ are isomorphic to the Riemann surface $\widetilde{\mathbb{C}}$ of the complex square root. This allows us to compare points in ${\widetilde{\mathbb{C}}}^{+}$ and ${\widetilde{\mathbb{C}}}^{-}$ by identifying these spaces with $\widetilde{\mathbb{C}}$. To achieve compatibility with [BO12, Section 6] and [HHP19] one finds that for both ${\widetilde{\mathbb{C}}}^{+}$ and $\widetilde{\mathbb{C}}^{-}$ one needs to choose $[\left\|\rho\right\|^{2},\infty)$ as the branch cut. For ${\widetilde{\mathbb{C}}}^{+}$ this corresponds to the “principal square root” branch cut of $\widetilde{\mathbb{C}}$ along $(-\infty,0]$, while for $\widetilde{\mathbb{C}}^{-}$ it corresponds to the “opposite” branch cut of $\widetilde{\mathbb{C}}$ along $[0,\infty)$. We define the physical sheet ~phys~subscriptsuperscript~$\widetilde{\mathbb{C}}^{-}_{\mathrm{phys}}\subset\widetilde{\mathbb{C}}^{-}$ over $\mathbb{C}\setminus[\left\|\rho\right\|^{2},\infty)$ as the sheet which is mapped bi-holomorphically onto $\mathcal{H}_{\mathrm{phys}}$ by the function with the -- sign in (4.6)
~$\widetilde{\mathbb{C}}^{-}_{\mathrm{phys}}\owns z\mapsto\sqrt{z-\left\|\rho\right\|^{2}}=:\zeta=:\zeta(z)\in\mathcal{H}_{\mathrm{phys}},$
providing us with the “physical parameter” $\zeta\in\mathcal{H}_{\mathrm{phys}}$.
Bunke and Olbrich, on the other hand, work with the parameter
$\lambda:=\lambda(z):=\sqrt{\left\|\rho\right\|^{2}-z}$
obtained by considering the $+$ sign in (4.6). For the conversion from $\lambda$ to $\zeta$ and vice-versa, there are the two possibilities $\lambda=\pm i\zeta$, which amount to choosing $\{\mp\mathrm{Re}\,\lambda>0\}\subset\mathbb{C}$ as the physical $\lambda$-half-plane corresponding to the physical $\zeta$-half-plane $\mathcal{H}_{\mathrm{phys}}$. In [BO12, p. 135] Bunke and Olbrich work with the right half-plane but their argument works just as well with the left half-plane. Indeed, there is no preferred choice of the physical sheet in [BO12, Section 6], so we can use their results regardless of which sign we choose in the relation between the parameters $\lambda$ and $\zeta$. The following choice will turn out to be convenient:
$\begin{split}\lambda&:=\lambda(\zeta):=-i\zeta\qquad\forall\;\zeta\in\mathcal{H}_{\mathrm{phys}},\\ \zeta&=\zeta(\lambda)=i\lambda,\quad\;\forall\lambda\in\mathbb{C},\;\mathrm{Re}\,\lambda>0.\end{split}$ (4.7)
@@ 4.3. Scattering matrices
Recall the critical exponent $\delta_{\Gamma}$ from (3.1) and note that if $\delta_{\Gamma}^{\mathrm{BO}}$ denotes the critical exponent from [BO00, Def. 2.2], then $\delta_{\Gamma}=\delta_{\Gamma}^{\mathrm{BO}}+\left\|\rho\right\|$.
@@
It is expected that the necessity of hypothesis $G\not=F_{4}^{-20}$ in Proposition 4.5 is an artefact of its proof. Note that the use of this result is the only reason why we had to restrict our attention to classical groups. We will address this issue elswhere.
Definition 4.6 (Scattering matrices for the globally symmetric space, c.f. [HHP19, §7]).
For $\zeta\in\mathcal{H}_{\mathrm{phys}}$ we define
$S_{\zeta}:=J_{i\zeta},$ (4.13)
the $\zeta$-scattering matrix for $X$.
Note that our $S_{\zeta}$ corresponds to $S_{-\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}}$ in [HHP19, §7]. This takes into account that our principal series representation parameter $\lambda$ corresponds to $-\lambda$ in [HHP19], while our resolvent $R_{\zeta}$ corresponds to $R_{-\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}}=R_{\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}}$ in [HHP19].

### v2

@@ 1. Introduction
The scattering poles mentioned in the title are the poles of the meromorphic families $S_{\lambda}$ and $S_{\lambda}^{\Gamma}$ of scattering matrices which in our situation can be constructed from the standard intertwining operators for the spherical principal series representations of $G$ using meromorphic families of extension operators constructed by Bunke and Olbrich in [BO00]. For technical reasons they had to exclude the case $G=F_{4}^{-20}$. This is the only reason why also we restrict our attention to the classical rank one spaces.
@@
The relation between quantum resonances and scattering poles for the global rank one symmetric spaces was described explicitly in [HHP19]. The quantum resonances are [^1] Note that the physical halfplane in [HHP19] is given by $\mathrm{Im}\,\zeta<0$, whereas we take $\mathrm{Im}\,\zeta>0$, which means that we need to flip the sign of $\zeta$ when formulating certain statements from [HHP19] in our setting. the zeros of the function $c(i\zeta)c(-i\zeta)$ with $\mathrm{Im}\,\zeta<0$, where $c$ is a variant of the Harish-Chandra $c$-function. The zeros of $c$ are known (see e.g. [HHP19, Rem. 5.1]). The main result [HHP19, Thm. 7.1] of that paper says that the scattering poles outside of $\mathcal{H}_{\mathrm{phys}}$ are precisely the quantum resonances (with equal and finite multiplicities), whereas the scattering poles which are not quantum resonances are the poles of the family of standard intertwining operators.
Earlier results are available (see [GZ97, BP02, Gu05]) for the case of asymptotically real hyperbolic manifolds. While they take the scattering matrices from geometric scattering theory and use microlocal analysis of manifolds with corners, the basic means of comparison between resonances and scattering poles is, just as in the harmonic analysis approach of [HHP19], the possibility to take boundary values of resolvent kernels yielding modifications of the scattering kernels. Borthwick’s book [Bo16] contains a very readable account of these results for the case of asymptotically hyperbolic surfaces.
That the results of Bunke and Olbrich could be used to extend the relation between quantum resonances and scattering poles to convex-cocompact locally symmetric spaces of rank one was suggested in [Hil24, Problem 6.5] based on the fact that in [BO12] Bunke and Olbrich studied the meromorphic extention of the resolvent kernels as well as the scattering poles for the classical locally symmetric spaces with negative curvature and convex-cocompact fundamental group.
@@
Our point of departure is [BO12, Lemma 11], where they show that the difference between the resolvent kernels on $X$ and the lift of the resolvent kernels on $\Gamma\backslash X$ has boundary values on $\partial X\times\partial X$ which depend meromorphically on the spectral parameters. A careful analysis of the proof of this lemma yields a meromorphic function $a$ relating the boundary values of the lifted resolvent kernels of $\Gamma\backslash X$ and the lifted kernels of the scattering matrices away from the diagonal. Our technical Lemma 5.3 (which may be of independent interest) shows that no extra singularities come from the diagonal and the results of [HHP19] allow to determine the function $a$. Thus we are lead to a generalization of the statement about the scattering poles being quantum resonances except for the exceptional spectral parameters for which the Poisson transform is singular so that no boundary value map is defined. More precisely, using the physical half-plane $\mathcal{H}^{\Gamma}_{\mathrm{phys}}$ depending on $\Gamma$, our main result can be formulated as follows, where $\alpha_{0}$ denotes the unique reduced root in a chosen positive system (see Theorem 5.4).
@@
Theorem. Outside the set $-\frac{i}{2}\left\|\alpha_{0}\right\|\mathbb{N}_{0}$, the scattering poles which are not in the physical half-plane $\mathcal{H}^{\Gamma}_{\mathrm{phys}}$ are precisely the quantum resonances. Moreover, if $\zeta_{0}\not\in-\frac{i}{2}\left\|\alpha_{0}\right\|\mathbb{N}_{0}$ is a simple pole of the quantum resolvent as well as the scattering matrix, then the boundary value map induces an isomorphism between the images of the two residue operators at $\zeta_{0}$.
In general, quantum resonances will not be simple poles of the resolvent, so our result on residues is incomplete. In the case of asymptotically hyperbolic manifolds complete results are available. While [BP02] achieve this by adding potentials to split spectral lines, [GZ97, Gu05] use Gohberg-Sigal normal forms (see [GS71]). It seems plausible to us that these techniques could also be employed in our situation, but we have not made a serious effort to do so as it would have lengthened the paper in a disproportionate way.
@@ Acknowledgements
Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 491392403 – TRR 358. BD has received further funding from the Deutsche Forschungsgemeinschaft through the Priority Program (SPP) 2026 “Geometry at Infinity”.
@@ 3. Resolvent kernels
$\mathcal{H}^{\Gamma}_{\mathrm{phys}}:=\begin{cases}\{\zeta\in\mathbb{C}\mid\mathrm{Im}\,\zeta>0\}&\text{for }0\leq\delta_{\Gamma}\leq\|\rho\|\\ \{\zeta\in\mathbb{C}\mid\mathrm{Im}\,\zeta>\delta_{\Gamma}-\|\rho\|\}&\text{for }\|\rho\|<\delta_{\Gamma}\leq 2\|\rho\|.\end{cases}$ (3.2)
[…]
@@
$c_{\Gamma}:=\delta_{\Gamma}^{2}+\max(0,\delta_{\Gamma}-\left\|\rho\right\|)^{2}\in[\delta_{\Gamma}^{2},2\delta_{\Gamma}^{2}]$ (3.4)
@@
$\forall\;\zeta\in W:\quad\big{|}((e,\gamma)^{\ast}r_{\zeta})(f)\big{|}$ $\leq C\left\|f\right\|_{\infty}e^{-a\sqrt{-\max(0,\delta_{\Gamma}-\left\|\rho\right\|)^{2}-\mathrm{Re}\,(\zeta^{2})}}$
@@ 4.1. Spherical principal series representations
$\sigma_{\lambda}:P=MAN\to\mathbb{C}\setminus\{0\},\quad man\mapsto a^{\rho+\lambda}$ (4.1)
@@
is an isomorphism of vector spaces. This allows us to consider $H^{\lambda,\infty}$ and $C^{\infty}(G/P;V(\sigma_{\lambda}))$ as $\lambda$-independent vector spaces. It also shows that $V(\sigma_{\lambda})$ is in fact a trivial line bundle but we will soon pass to its $\Gamma$-quotient, which no longer needs to be trivial.
@@
over the smooth manifold $\Gamma\backslash\Omega$. The vector space $C^{\infty}(\Gamma\backslash\Omega;V_{\Gamma\backslash\Omega}(\sigma_{\lambda}))$ corresponds canonically to the space of $\Gamma$-invariant smooth sections of $V(\sigma_{\lambda})|_{\Omega}$, which in turn corresponds canonically to a closed subspace of left-$\Gamma$-invariant functions in $H^{\lambda,\infty}$. Thus, the space $C^{\infty}(\Gamma\backslash\Omega;V_{\Gamma\backslash\Omega}(\sigma_{\lambda}))$ inherits a Fréchet topology from $H^{\lambda,\infty}$. In particular, we can consider $C^{\infty}(\Gamma\backslash\Omega;V_{\Gamma\backslash\Omega}(\sigma_{\lambda}))$ as a $\lambda$-independent Fréchet space, too.
@@ 4.2. Comparing conventions
Remark 4.2 (Signs and left/right regular representation).
Our definitions (4.1), (4.2) differ from the ones in [BO00, Section 3], [Olb94, § 2.2] and [HHP19, § 2] by the sign of $\lambda$. We chose this sign convention to be consistent with [Kna86, VII, §1]. This is also consistent with the sign convention in [Vog08, Chap. 3]. In all these references the $G$-action is the left regular action. On the other hand, [Wal92, § 10.1] and [KS71] use the right regular version of the principal series representations. Note that the map $f\mapsto\check{f}$ with $\check{f}(g):=f(g^{-1})$ intertwines left and right regular actions. Up to this intertwiner the sign conventions in [Wal92, § 10.1] and [KS71] agree with the sign conventions in [Kna86, VII, §1]. Finally, we point out that our topologies on $C^{\infty}(G/P;V(\sigma_{\lambda}))$ and $C^{\infty}(\Gamma\backslash\Omega;V_{\Gamma\backslash\Omega}(\sigma_{\lambda}))$ are equivalent to those of [BO00, p. 89].
Remark 4.3 (Norms and identifications).
@@
We use the convention from [BO12, § 6.1] that the complex vector space ${\bf\mathfrak{a}}^{\ast}_{\mathbb{C}}$ is identified with $\mathbb{C}$ by complex linear extension of the unique order-preserving isometric isomorphism ${\bf\mathfrak{a}}^{\ast}\cong\mathbb{R}$. This means that $\rho\in{\bf\mathfrak{a}}^{\ast}$ is identified with $\left\|\rho\right\|\in\mathbb{R}$ (which accordingly is written simply as $\rho$ in [BO12, § 6]), and a general a number $z\in\mathbb{C}$ corresponds to the element
$\frac{z}{\left\|\rho\right\|}\rho\in{\bf\mathfrak{a}}^{\ast}_{\mathbb{C}}.$ (4.6)
This identification differs slightly from that in [HHP19], where $\rho\in{\bf\mathfrak{a}}^{\ast}$ gets identified with $\frac{\left\|\rho\right\|}{\left\|\alpha_{0}\right\|}\in\mathbb{R}$, $\alpha_{0}\in\Sigma^{+}\subset{\bf\mathfrak{a}}^{\ast}$ being the unique reduced positive root, and consequently a general number $z\in\mathbb{C}$ corresponds to $\left\|\alpha_{0}\right\|$ times the element (4.6).
Remark 4.4 (Branched covers, physical sheets and half-planes).
Our choice of the physical half-plane $\mathcal{H}_{\mathrm{phys}}=\{\zeta\in\mathbb{C}\,|\,\mathrm{Im}\,\zeta>0\}$ is compatible with [HHP19] when taking into account the sign change mentioned in Remark 4.2. On the other hand, Bunke and Olbrich work in [BO12, p. 135] with the right half-plane $\mathcal{H}_{\mathrm{phys,r}}^{\mathrm{BO}}:=\{\lambda\in\mathbb{C}\,|\,\mathrm{Re}\,\lambda>0\}$ but their argument works just as well with the left half-plane $\mathcal{H}_{\mathrm{phys,l}}^{\mathrm{BO}}:=\{\lambda\in\mathbb{C}\,|\,\mathrm{Re}\,\lambda<0\}$.
Let ${\widetilde{\mathbb{C}}}^{\pm}$ be the branched covers of $\mathbb{C}$ on which the functions
~±${\widetilde{\mathbb{C}}}^{\pm}\owns z\mapsto\sqrt{\pm(\left\|\rho\right\|^{2}-z)}=:\zeta_{\pm}(z)\in\mathbb{C}$ (4.7)
are holomorphic, respectively. In [BO12, p. 135] Bunke and Olbrich work with the function $\zeta_{+}$, while Hansen, Hilgert and Parthasarathy work with $\zeta_{-}$ in [HHP19]. We call ~$\widetilde{\mathbb{C}}^{-}_{\mathrm{phys}}:=\zeta_{-}^{-1}(\mathcal{H}_{\mathrm{phys}})\subset\widetilde{\mathbb{C}}^{-}$ the physical sheet of $\widetilde{\mathbb{C}}^{-}$ and ~$\widetilde{\mathbb{C}}^{+}_{\mathrm{phys,r}}:=\zeta_{+}^{-1}(\mathcal{H}_{\mathrm{phys,r}}^{\mathrm{BO}}),\widetilde{\mathbb{C}}^{+}_{\mathrm{phys,l}}:=\zeta_{+}^{-1}(\mathcal{H}_{\mathrm{phys,l}}^{\mathrm{BO}})\subset\widetilde{\mathbb{C}}^{+}$ the positive/negative physical sheets of $\widetilde{\mathbb{C}}^{+}$, respectively. We emphasize that there is no preferred choice of the physical sheet in [BO12, Section 6], so we can use their results for parameters in $\mathcal{H}_{\mathrm{phys,r}}^{\mathrm{BO}}$ as well as in $\mathcal{H}_{\mathrm{phys,l}}^{\mathrm{BO}}$.
As before, we will denote the complex numbers in our physical half-plane by $\zeta$, while we will write $\lambda$ for the numbers in $\mathcal{H}_{\mathrm{phys,r}}^{\mathrm{BO}}$ and $\mathcal{H}_{\mathrm{phys,l}}^{\mathrm{BO}}$. We have obvious isomorphisms $\mathcal{H}_{\mathrm{phys}}\cong\mathcal{H}_{\mathrm{phys,r}}^{\mathrm{BO}}$, $\mathcal{H}_{\mathrm{phys}}\cong\mathcal{H}_{\mathrm{phys,l}}^{\mathrm{BO}}$ given by
$\mathcal{H}_{\mathrm{phys}}\owns\zeta\mapsto\lambda=\begin{cases}-i\zeta&\in\mathcal{H}_{\mathrm{phys,r}}^{\mathrm{BO}},\\ i\zeta&\in\mathcal{H}_{\mathrm{phys,l}}^{\mathrm{BO}}.\end{cases}$ (4.8)
@@ 4.3. Scattering matrices
Recall the critical exponent $\delta_{\Gamma}$ from (LABEL:eq:criticalexponent) and note that if $\delta_{\Gamma}^{\mathrm{BO}}$ denotes the critical exponent from [BO00, Def. 2.2], then $\delta_{\Gamma}=\delta_{\Gamma}^{\mathrm{BO}}+\left\|\rho\right\|$.
@@
We expect that the necessity of the hypothesis $G\not=F_{4}^{-20}$ in Proposition 4.6 is an artefact of its proof. Recall that the use of this result is the only reason why we had to restrict our attention to classical groups. We will address this issue elsewhere.
Definition 4.7 (Scattering matrices for the locally and globally symmetric space, c.f. [HHP19, §7]).
We call the two meromorphic operator families
$\mathbf{S}_{\zeta}$ $:=J_{i\zeta},$
$\mathbf{S}_{\zeta}^{\Gamma}$ $:=S^{\Gamma}_{i\zeta}$
the $\zeta$-scattering matrices for $X$.
Note that our $\mathbf{S}_{\zeta}$ corresponds to $S_{-\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}}$ in [HHP19, §7]. This takes into account that our principal series representation parameter $\lambda$ corresponds to $-\lambda$ in [HHP19], while our resolvent $R_{\zeta}$ corresponds to $R_{-\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}}=R_{\frac{\zeta\alpha_{0}}{\left\|\alpha_{0}\right\|}}$ in [HHP19].